OpenStudy (anonymous):

how long will it take $3000 to grow to $20000 if compounded monthly at 7%

OpenStudy (solomonzelman):

\(\huge\color{blue}{ A=P(1+\frac{r}{n})^{nt} }\) P = principal amount (the initial amount you borrow or deposit) r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year plug your numbers in. \(\huge\color{blue}{ 20,000=3,000(1+\frac{7}{12})^{(12)t} }\) need more help solving for "t" - (the number of years) ?

OpenStudy (solomonzelman):

any questions, PLEASE ASK, don't leave. (I'll be back when you reply)

OpenStudy (anonymous):

t...?

OpenStudy (anonymous):

@ SolomanZelman how did you put value of N.....?